Emily takes a trip, driving with a constant velocity of 81.5 km/h to the north except for a 22 min rest stop. If Emily's average velocity is 72.8 km/h to the north, how long does the trip take?
1
Expert's answer
2012-10-11T08:38:00-0400
Let's denote the driving time by T. Then the travelled distance can be expressed in a such way:

D = 72.8[km/h]*(T[h]+22[m]) = 72.8[km/h]*(T[h]+11/30[h]).

From the other side,

D = 81.5[km/h]*T.

So,

72.8[km/h]*(T[h]+11/30[h]) = 81.5[km/h]*T ==>

72.8*T + 72.8*11/30 = 81.5*T ==>

800.8/30 = 8.7*T ==>

T ≈ 3.09[h].

So, the trip took about (3.09 + 11/30) ≈ 3.46 hours.